OpenStudy (anonymous):

Find two linearly independent solutions using the recursive relation (power series) of y''+xy'-3y=0

OpenStudy (anonymous):

is the equation \[y^2+xy-3y=0\]

OpenStudy (kainui):

\[y=\sum_{n=0}^{\infty}a_nx^n\] Try using this and show me how far you get.

OpenStudy (anonymous):

I got the recursive relation as \[c _{n+2}=\frac{ 3-n }{ (n+2)(n+1) }c_{n}\]